OpenStudy (anonymous):

Letf(x,y,z)=x+ln(2y+3z^2) compute ∂2f/∂y∂x

OpenStudy (anonymous):

You can differentiate in any order, so do the easiest first: df/dx = 1 + 0. So df/dxdy = 0. done

OpenStudy (anonymous):

i thought df/dy was2/2y+3z^2?

OpenStudy (anonymous):

It is, but you can differentiate with respect to x first. You can check it by differentiating df/dy with respect to x. You still get zero.

OpenStudy (anonymous):

so then Letf(x,y,z)=x+ln(2y+3z^2) compute ∂^3f/∂y∂x∂z is 0 to?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

This website will differentiate stuff like this http://www.wolframalpha.com input it like this http://www.wolframalpha.com/input/?i=differentiate+x%2Bln%282y%2B3z^2%29+with+respect+to+x